# Recent Advances in Discrete Speech Tokens: A Review

### v2

@@ III-D3 Challenges
The GRL technique for disentanglement inherently carries the risk of a more complex optimization trajectory. Additionally, some disentanglement methods require supervised data [28], which imposes a significant constraint. Due to the intricate nature of speech informatics, current efforts are still suboptimal compared to semantic tokens, particularly in terms of VC performance [29].
@@ IV Speech Tokenization Methods: Semantic Tokens
Semantic tokens refer to discrete speech representations from discriminative or self-supervised learning (SSL) models. While we use the term semantic tokens to maintain consistency with prior works, some researchers recently argue that SSL features are more accurately described as phonetic than semantic [113] in nature. Hence to clarify, in this review, semantic tokens should be more accurately defined as the complementary set of acoustic tokens, such that they are not primarily aimed at reconstruction purposes. In practice, the vast majority of these tokens are designed for discriminative tasks and are believed to have a strong correlation with phonetic and semantic information [114, 8, 115, 116].
IV-A Semantic Tokens from General-Purpose SSL
@@ IV-A1 Motivation
Speech SSL models have consistently outperformed many traditional methods in various speech tasks [117, 8]. Their potential has been extensively mined in discriminative tasks such as automatic speech recognition (ASR) [118, 31, 33, 119], automatic speaker verification (ASV) [34, 120, 121], speech emotion recognition (SER) [122, 34, 123, 124] and speech translation (ST) [125, 126, 127]. Discretized SSL tokens are initially favored for reducing computation costs and improving robustness against irrelevant information for ASR [128]. As language models have gained increasing attention, these SSL tokens have been further explored in generative tasks such as TTS [129, 15, 130] and SLM [4, 5, 131]. This is because they can be considered high-level abstractions of speech semantics that are largely independent of acoustic details.
Figure 7: Representatives in different kinds of semantic tokens. “Q.” denotes quantizer, which can be optional (dotted line).
@@ IV-A2 Approaches
  • External quantization, like clustering or training a VQ-VAE. This refers to extracting continuous embeddings from a certain layer or multiple layers in a pretrained SSL model, and performing quantization manually. For example, a common semantic token is the HuBERT+kmeans units, where k-means clustering is performed on a HuBERT Transformer layer with a portion of training data [4, 132]. It is also feasible to perform clustering on multiple layers [133, 134], or train a VQ-VAE on the SSL hidden embeddings [135, 136].
@@
  • When an SSL model contains an inner quantizer that is trained together with other network modules, its outputs can also be regarded as semantic tokens. Many SSL models involve quantizers to produce targets for their training objectives [31, 32, 137, 138]. This approach provides an efficient and effective way of extracting semantic tokens.
@@
Note that for SSL models with an inner quantizer, it is still practical to perform external quantization on its continuous embeddings, like wav2vec 2.0 [32]. However, these two methods – internal and external quantization – may result in different patterns of information exhibition, which we will investigate in Section V.
@@
For general-purpose SSL models, there are different designs on the pretext task [8]. Table II provides a high-level summary of well-known semantic tokens.
Contrastive
@@
This type of speech SSL models aims to learn representations by distinguishing a target sample (positives) from distractors (negatives) given an anchor [8]. They minimize the latent space similarity of negative pairs and maximize that of the positive pairs. For semantic tokens, vq-wav2vec [31] and wav2vec 2.0 [32] are two representative contrastive SSL models. They involve a quantizer to produce localized features that is contrastively compared to contextualized continuous features. Vq-wav2vec [31] uses pure CNN blocks while wav2vec 2.0 [32] adopts a Transformer for stronger capacity. Both use GVQ quantizers with two groups to expand the VQ space. Wav2vec 2.0 has also been extended to massively multilingual versions [139, 127, 140].
Predictive
@@
This type of speech SSL models incorporates an external target for prediction, either from signal processing features or another teacher network. A popular line of work is HuBERT [33]. It takes raw waveforms as inputs, applies random masks on the hidden representations before Transformer contextual blocks, and then predicts k-means quantized targets from MFCC or another HuBERT teacher. WavLM [34] augments HuBERT by additional speaker and noise perturbations to achieve superior performance in more paralinguistic-related tasks. There are no inner quantizers in both models, so external quantization like k-means clustering is necessary to obtain semantic tokens. BEST-RQ [137] changes the prediction target to the output of a random projection quantizer. The next-token prediction criterion from language models (LMs) have also been adopted into speech SSL [141, 142], either with or without a pretrained text LM. This method emphasizes the autoregressive prediction property of learned tokens that may be better suited for the LM use case.
@@ IV-A3 Challenges
Firstly, SSL models typically require large amount of data to train, as indicated in Table II. For SSL models without a built-in quantizer during pretraining, k-means clustering is a prevalent approach to obtain discrete units. However, given that most SSL models operate in high-dimensional spaces (e.g., with 768 or 1024 dimensions), the space and time complexity of k-means clustering are substantial. The clustering results can sometimes be unreliable due to the curse of dimensionality in Euclidean space. Moreover, it is often reported, and will also be shown by experiments in Section V, that discretized SSL units lose much acoustic details after quantization [143, 115, 144]. Different clustering settings, such as the chosen layer and vocabulary size, can lead to different outcomes within a single model. Finally, since most SSL models utilize Transformer blocks, their causality and streaming ability are compromised.
IV-B Semantic Tokens from Perturbation-Invariant SSL
IV-B1 Motivation
As SSL tokens feature semantic or phonetic information, a major concern is to improve the resistance against perturbations in the input signal. This kind of invariance includes noise and speaker aspects that don’t affect the contents of speech. Noise invariance refers to the invariance against signal augmentations such as additive noise, reverberations, etc. Speaker invariance aims to remove speaker information, similar to speaker-disentangled acoustic tokens. In the training process, perturbations are often explicitly introduced in these perturbation-invariant SSL models. The original and perturbed view of an utterance are both fed to the same network (or teacher and student networks), and an external loss to reduce the impact of perturbation is applied. The middle part of Fig.7 depicts a typical perturbation-invariant SSL model.
IV-B2 Approaches
Perturbations
The perturbations can either be designed to augment the acoustics or alter the speaker timbre, depending on the objective of invariance. These perturbations usually preserve temporal alignments, meaning that the perturbed utterance and the original one are strictly synchronized. For noise-invariant SSL tokens, basic signal variations like time stretching, pitch shifting, additive noise, random replacing, reverberation, and SpecAugment [145] are commonly applied [39, 146, 38, 36]. Typical speaker timbre perturbations include formant and pitch scaling as well as random equalization [35, 37, 147]. In contrast, random time stretching is applied as speaker perturbation in [148], which alters the tempo in each random segment.
Contrastive-based Methods
Contrastive loss is a common method to obtain perturbation-invariant representations. In this context, the contrastive loss is a modified version of that used in wav2vec 2.0 [32]. Given two embedding sequences derived from the original and perturbed utterances, assuming the perturbation preserves frame-wise alignment, the positive sample of an anchor is taken from the same position in the other utterance. This is because the content remains unchanged by the perturbation, thus the same position of two representation sequences should encode the same information. In noise-invariant models [36, 146], negative samples are selected from the other utterance relative to the anchor. However, in speaker-invariant models [35, 148], negative samples are selected from the same utterance as the anchor. Specifically, in [148], soft attention pooling is applied to create equal-length representation sequences from two utterances with different durations. This approach forces SSL models to ignore acoustic differences and focus solely on the unperturbed content.
Distribution-based Methods
Another method to achieve invariance is to minimize some distance metrics between the representations extracted from the original and perturbed utterances. In existing perturbation-invariant SSL models, this is typically accomplished using a cross-entropy loss between the underlying distributions in the VQ module of the SSL model. NAST [38] trains a Gumbel-based VQ-VAE on HuBERT features and enforces similarity between the Gumbel distributions Eq.(3) derived from the original and augmented utterances. Spin [37] and DC-Spin [147] explore a speaker-invariant clustering algorithm for HuBERT features. Similar to NAST [38], Spin employs a cross-entropy loss to ensure that the distributions over codebook entries are similar between the original and perturbed utterances. This distribution-based approach forces the same content to be quantized to the same index regardless of acoustic conditions.
CTC-based Methods
Noise invariance can also be achieved like an ASR task with perturbed speech inputs. As semantic tokens from SSL models are highly content-related, these tokens extracted from the original clean utterance can serve as some pseudo-label for a perturbed view. In [39], a connectionist temporal classification (CTC) [149] loss is calculated between quantized tokens from the augmented signal and a pretrained HuBERT+kmeans pseudo-labels from the clean signal. This pushes the quantized tokens to have the same phonetic structure with the pseudo-labels.
IV-B3 Challenges
@@
While noise and speaker-invariance have emerged as promising approaches in semantic tokens, they currently rely on content-preserving perturbations that are typically hand-crafted. Most existing methods have only been evaluated on small-scale data and models. It also remains unclear how these methods will generally benefit generative tasks such as speech generation and spoken language modeling.
TABLE II: A high-level summary of famous semantic speech tokens. Notations follow Table.I. Symbol ‘/’ denotes different versions. “Inner Quantizer” refers to whether the model has a quantizer, or external quantization (e.g. clustering) must be performed. $F$ denotes frame rate. In case there are inner quantizers, $Q,V$ denote number of quantizers and vocabulary size for each quantizer, respectively. “NR.” means not reported.
General-purpose self-supervised learning (SSL) models
vq-wav2vec [31] Contrastive 0.96k 100 GVQ, $Q=2,V=320$
wav2vec 2.0 [32] Contrastive 60k 50 GVQ, $Q=2,V=320$
XLSR-53 [139] Contrastive 50k 50 GVQ, $Q=2,V=320$
HuBERT [33] Predictive 60k 50 No
WavLM [34] Predictive 94k 50 No
BEST-RQ [137] Predictive 60k 25 No
w2v-BERT [150] Predictive+Contrastive 60k 50 VQ, $Q=1,V=1024$
w2v-BERT 2.0 [107] Predictive+Contrastive 4500k 50 GVQ, $Q=2,V=320$
DinoSR [151] Predictive 0.96k 50 VQ, $Q=8,V=256$
NEST-RQ [142] Predictive 300k 25 No
LAST [141] Predictive NR. 50 VQ, $Q=1,V=500$
SSL models with perturbation-invariance
Gat et al. [39] Noise Invariance 0.10k 50 VQ, $G=1,V=50$-$500$
ContentVec [35] Speaker Invariance 0.96k 50 No
SPIRAL [36] Noise Invariance 60k 12.5Hz No
CCC-wav2vec 2.0 [146] Noise Invariance 0.36k 50 GVQ, $G=2,V=320$
Spin [37] Speaker Invariance 0.10k 50 VQ, $Q=1,V=128$-$2048$
NAST [38] Noise Invariance 0.96k 50 VQ, $Q=1,V=50$-$200$
DC-Spin [147] Speaker Invariance 0.96k 50 VQ, $Q=2,V=(50$-500)500)500 )+$4096$
Supervised models
$\mathcal{S}^{3}$ Tokenizer [41] Supervised ASR 172k 25 / 50 VQ, $Q=1,V=4096$
Zeng et al. [152] Supervised ASR 90k 12.5 VQ, $Q=1,Q=16384$
Du et al. (CosyVoice 2) [153] Supervised ASR 200k 12.5 FSQ, $Q=8,V=3$
IV-C Semantic Tokens from Supervised Models
@@
As representing semantic or phonetic information is the major purpose of semantic tokens, a more direct way to achieve this is through supervised learning. A famous example shown at the bottom of Fig.7 is the $\mathcal{S}^{3}$ Tokenizer from CosyVoice [41]. It places a single-codebook VQ layer between two Transformer encoder modules and optimizes the network through an ASR loss similar to Whisper [40]. The same method is adopted in [152, 154] where the frame rate is further reduced to 12.5Hz. CosyVoice 2 [153] improves $\mathcal{S}^{3}$ Tokenizer by replacing plain VQ with FSQ for better codebook utilization. Note that in this kind of supervised semantic tokens, it is the output of the VQ layer that serves as tokens. This allows for more preservation of paralinguistic information than directly transcribing speech into text. These supervised tokenizers are trained on massive paired speech-text data, and have demonstrated rich speech content understanding capabilities [41, 155].
@@
However, training these models is highly costly due to the heavy data demands. Training with only the ASR task may still result in the loss of some prosody information. Although [153] has demonstrated that its supervised tokenizer trained on Chinese and English can also work in Japanese and Korean, it remains unclear how well these supervised tokenizers generalize to more unseen languages.
IV-D Length Reduction by Deduplication and Acoustic BPE
@@
In most cases, the frame rate of discrete speech tokens ranges from 25 to 100Hz. This leads to a huge discrepancy in lengths between speech representations and the underlying text modality. This discrepancy has been a critical issue in building decoder-only TTS and other LM-based speech generation tasks, since longer sequences result in harder training and more unstable inference. Therefore, length reduction techniques have been proposed to address this issue. These methods are inspired by language processing techniques and are thus more closely related to semantic tokens. Note that although these length reduction methods are universal across token types, they are less frequently applied to acoustic tokens. This is because acoustic tokens usually involve multiple VQ streams that complicate token-level operations.
@@
A common approach to reduce token sequence lengths is deduplication [128, 156], i.e. removing the repeated consecutive tokens in a sequence. Since the encoded continuous features are often close in consecutive frames where the speech dynamics do not change rapidly, they are likely to be quantized to the same unit. Therefore, removing these redundant tokens can yield a more phonetic representation. When the deduplicated tokens are used for generative modeling, a unit-to-speech model (similar to TTS) should be employed to upsample the tokens and convert them back to acoustic signals [4].
@@
Another popular approach to reducing the length of speech token sequences is acoustic byte-pair encoding (BPE) [^4] The term “acoustic” here is used to distinguish it from traditional BPE applied to text tokens, rather than referring to “acoustic tokens”. or so-called subword modeling [157, 158, 128, 159, 160]. Similar to text BPE [161], acoustic BPE iteratively merges the two most frequent consecutive tokens and adds the merged token to the vocabulary. After training on a corpus, a deterministic BPE mapping is established between original token combinations and the new vocabulary. This mapping enables a lossless compression algorithm, allowing tokens to be perfectly reconstructed after BPE decoding. This operation can identify certain morphological patterns in token sequences, and offers a powerful way to remove redundant tokens. In practice, acoustic BPEs on HuBERT semantic tokens has demonstrated significant speed and performance gains in ASR [128, 156], spoken language modeling [159, 160] and TTS [162, 130].
@@
Although deduplication is a simple and training-free method, acoustic BPE offers several unique advantages over it. First, acoustic BPE can identify redundant patterns that are not simply repetitions, whereas deduplication only removes exact duplicates. Additionally, deduplication discards the duration information of every token in the resulting sequence. This could be problematic for downstream tasks, as important rhythmic information may reside in the repetitions of tokens. In contrast, acoustic BPE preserves duration information by encoding repetitions of varying lengths into distinct new tokens. Furthermore, acoustic BPE is flexible in terms of target vocabulary size, which can be adjusted based on the desired length reduction ratio and downstream performance.
@@
We visualize the length reduction effect of BPE on different speech tokens in Fig.8. In addition to semantic tokens from various models and different k-means clusters in HuBERT, we also experiment with acoustic tokens. For acoustic tokens with multiple codebooks, we apply BPE only to the first quantizer, in accordance with the current speech generation paradigm [6]. From Fig.8, it is evident that different types of tokens exhibit very distinct patterns. Semantic tokens generally show significant length reduction when applying BPE, especially for HuBERT models with fewer k-means clusters. For single-codebook acoustic tokens, speaker-decoupled LSCodec tokens shows more reduction than general-purpose WavTokenizer and BigCodec. For a single RVQ layer among multiple-codebook acoustic tokens, the reduction effect is also significant. These findings suggest that the effect of BPE is negatively correlated with the information density in the speech tokens: the less information, the more length reduction achieved by BPE.
IV-E Variable Frame Rate Tokens and Unit Discovery

### v3

@@ III-D3 Challenges
The GRL technique for disentanglement inherently carries the risk of a more complex optimization trajectory. Additionally, some disentanglement methods require supervised data [facodec], which imposes a significant constraint. Due to the intricate nature of speech informatics, current efforts are still suboptimal compared to semantic tokens, particularly in terms of VC performance [guo2024lscodec].
@@ IV Speech Tokenization Methods: Semantic Tokens
Semantic tokens refer to discrete speech representations from discriminative or self-supervised learning (SSL) models. While we use the term semantic tokens to maintain consistency with prior works, some researchers recently argue that SSL features are more accurately described as phonetic than semantic [choi24b_interspeech] in nature. Hence to clarify, in this review, semantic tokens should be more accurately defined as the complementary set of acoustic tokens, such that they are not primarily aimed at reconstruction purposes. In practice, the vast majority of these tokens are designed for discriminative tasks and are believed to have a strong correlation with phonetic and semantic information [wells22_interspeech, mohamed2022self, sicherman2023analysing, yeh2024estimating].
IV-A Semantic Tokens from SSL Models
@@ IV-A1 Motivation
When fine-tuned, speech SSL models have shown strong performance on various tasks, often surpassing traditional methods [superb, mohamed2022self]. Their potential has been extensively mined in discriminative tasks such as automatic speech recognition (ASR) [wav2vec, vq-wav2vec, hsu2021hubert, zhang2020pushing], automatic speaker verification (ASV) [chen2022wavlm, jung2024espnet, miara24_interspeech], speech emotion recognition (SER) [morais2022speech, chen2022wavlm, MADANIAN2023200266, ma-etal-2024-emotion2vec] and speech translation (ST) [wu20g_interspeech, nguyen20_interspeech, babu22_interspeech]. Discretized SSL tokens are initially favored for reducing computation costs and improving robustness against irrelevant information for ASR [chang23b_interspeech]. Driven by the success of language models, these SSL tokens have been further explored in generative tasks such as TTS [VQTTS, kharitonov2023speak, vectokspeech] and SLM [lakhotia2021generative, borsos2023audiolm, hassid2024textually]. This is because they can be considered high-level abstractions of speech semantics that are largely independent of acoustic details.
Figure 7: Representatives in different kinds of semantic tokens. Dotted box means the module is optional.
@@ IV-A2 Approaches
  • External quantization, like clustering or training a VQ-VAE. This refers to extracting continuous embeddings from a certain layer or multiple layers in a pretrained SSL model, and performing quantization manually. For example, a common semantic token is the HuBERT+kmeans units, where k-means clustering is performed on a HuBERT Transformer layer with a portion of training data [lakhotia2021generative, kharitonov-etal-2022-text]. It is also feasible to perform clustering on multiple layers [shi24h_interspeech, mousavi2024should], or train a VQ-VAE on the SSL hidden embeddings [huang2023repcodec, wang2024maskgct].
@@
  • Internal quantization: when an SSL model contains an inner quantizer that is trained together with other network modules, its outputs can also be regarded as semantic tokens. Many SSL models involve quantizers to produce targets for their training objectives [vq-wav2vec, baevski2020wav2vec, chiu2022self, zhu2025muq]. This approach provides an efficient and effective way of extracting semantic tokens.
@@
Note that for SSL models with an inner quantizer, it is still practical to perform external quantization on its continuous embeddings, like wav2vec 2.0 [baevski2020wav2vec]. However, these two methods – internal and external quantization – may result in different patterns of information exhibition, which we will investigate in Section VI.
@@
For general-purpose SSL models, there are different designs on the pretext task [mohamed2022self]. Table III provides a summary of well-known semantic tokens.
Contrastive models
@@
This type of speech SSL models aims to learn representations by distinguishing a target sample (positives) from distractors (negatives) given an anchor [mohamed2022self]. They minimize the latent space similarity of negative pairs and maximize that of the positive pairs. For semantic tokens, vq-wav2vec [vq-wav2vec] and wav2vec 2.0 [baevski2020wav2vec] are two representative contrastive SSL models. They involve a quantizer to produce localized features that is contrastively compared to contextualized continuous features. Vq-wav2vec [vq-wav2vec] uses pure CNN blocks while wav2vec 2.0 [baevski2020wav2vec] adopts a Transformer for stronger modeling capacity. Both use GVQ quantizers to expand the VQ space. Wav2vec 2.0 has also been extended to massively multilingual versions [conneau21_interspeech, babu22_interspeech, pratap2024scaling].
Predictive models
@@
This type of speech SSL models incorporates an external target for prediction, either from signal processing features or another teacher network. A popular line of work is HuBERT [hsu2021hubert]. It takes raw waveforms as inputs, applies random masks on the hidden representations before Transformer contextual blocks, and then predicts k-means quantized targets from MFCC or another HuBERT teacher. WavLM [chen2022wavlm] improves HuBERT by additional speaker and noise augmentations to achieve superior performance in more paralinguistic-related tasks. There are no inner quantizers in both models, so external quantization like k-means clustering is necessary to obtain semantic tokens. BEST-RQ [chiu2022self] changes the prediction target to the output of a random projection quantizer. The next-token prediction criterion from language models (LMs) have also been adopted into speech SSL [turetzky2024last, han2024nest], either with or without a pretrained text LM. This method emphasizes the autoregressive prediction property of learned tokens that may be better suited for the LM use case.
Perturbation-invariant models
As SSL tokens feature semantic or phonetic information, a major concern is to improve the resistance against perturbations in the input signal. This invariance includes noise and speaker aspects that don’t affect the contents of speech. Specifically, speaker-invariant SSL tokens aims to remove speaker information, similar to speaker-disentangled acoustic tokens. In the training process of perturbation-invariant SSL models, noise [gat2023augmentation, ccc-wav2vec2.0, messica2024nast, huang2022spiral] or speaker [qian2022contentvec, chang23_interspeech, chang2024dc, hwang2024removing] augmentations are often explicitly introduced. Special training losses are then incorporated, like contrastive losses [huang2022spiral, ccc-wav2vec2.0, qian2022contentvec, hwang2024removing] that distinguish the same spoken content among perturbed distractors, and distribution-based metrics [chang23_interspeech, messica2024nast, chang2024dc] that minimize the distance of latent features caused by perturbations.
@@ IV-A3 Challenges
Firstly, SSL models typically require large amount of data to train, as indicated in Table III. For SSL models without a built-in quantizer during pretraining, k-means clustering is a prevalent approach to obtain discrete units. However, given that most SSL models operate in high-dimensional spaces (e.g., with 768 or 1024 dimensions), the space and time complexity of k-means clustering are substantial. The clustering results can sometimes be unreliable due to the curse of dimensionality in Euclidean space. Moreover, it is often reported, and will also be shown by experiments in Section VI, that discretized SSL units lose much acoustic details after quantization [polyak21, sicherman2023analysing, mousavi2024dasb]. Different clustering settings, such as the chosen layer and vocabulary size, can lead to different outcomes within a single model. Finally, since most SSL models utilize Transformer blocks, their causality and streaming ability are compromised.
@@
While perturbation-invariant SSL models have emerged as promising approaches for semantic tokens, they currently rely on content-preserving augmentations that are typically hand-crafted. Most methods in this type have only been evaluated on small-scale data and models. It also remains unclear how these methods will generally benefit generative tasks such as speech generation and spoken language modeling.
IV-B Semantic Tokens from Supervised Models
@@
As representing semantic or phonetic information is the major purpose of semantic tokens, a more direct way to achieve this is through supervised learning. Supervised semantic tokenizers are typically trained on the ASR task. A famous example shown at the bottom of Fig.7 is the $\mathcal{S}^{3}$ Tokenizer from CosyVoice [du2024cosyvoice]. It places a single-codebook VQ layer between two Transformer encoder modules and optimizes the network using a cross-entropy ASR loss, similar to Whisper [whisper]. The same method is adopted in [zeng2024scaling, zeng2024glm] where the frame rate is further reduced to 12.5Hz. CosyVoice 2 [cosyvoice2] improves $\mathcal{S}^{3}$ Tokenizer by replacing plain VQ with FSQ for better codebook utilization. Note that in this kind of supervised semantic tokens, it is the output of the VQ layer that serves as tokens. This allows for more preservation of paralinguistic information than directly transcribing speech into text. CosyVoice 3 [du2025cosyvoice3] extends supervised tokens to more tasks involving language, emotion, speaker and audio analysis. These supervised tokenizers are trained on massive paired speech-text data, and have demonstrated rich speech content understanding capabilities [du2024cosyvoice, fang2024llamaomni].
@@
However, training these models is highly costly due to the heavy data demands. Training with only the ASR task may still result in the loss of some prosody information, and training with multiple tasks poses a higher demand for data and task balancing. Although [cosyvoice2] has demonstrated that its supervised tokenizer trained on Chinese and English can also work in Japanese and Korean, it remains unclear how well these supervised tokenizers generalize to more unseen languages.
IV-C Speech Token Vocoders
Acoustic tokens are naturally coupled with a decoder that outputs waveforms or spectrograms given tokens, but semantic tokens do not inherently require a vocoder, particularly for speech understanding tasks. However, when semantic tokens are used for speech generation, a speech token vocoder (also known as speech resynthesis model) becomes necessary. Unlike traditional spectrogram-based vocoders [kong2020hifigan], speech token vocoders usually need to compensate the loss of acoustic details in semantic tokens by introducing additional inputs.
Polyak et al. [polyak21] first explores speech resynthesis from discrete speech units by a HifiGAN [kong2020hifigan] augmented with discretized pitch units and speaker embedding inputs. The vec2wav vocoder in VQTTS [VQTTS] improves this vocoder by a Conformer [conformer] frontend module before HifiGAN generator. Later, CTX-vec2wav [du2024unicats] proposes a position-agnostic cross-attention mechanism that effectively integrates timbre information from surrounding acoustic contexts without the need for pretrained speaker embeddings. This makes it more timbre-controllable and suitable for zero-shot TTS and VC [li2024sef]. Upon it, vec2wav 2.0 [guo2024vec2wav] introduces SSL timbre features and adaptive activations to improve timbre controllability, and reports competitive VC performance.
It is also feasible to apply diffusion or flow matching algorithms in token vocoders [tortoise, seedtts, du2024cosyvoice]. There, the discrete tokens are treated as a condition for diffusion or flow matching to generate mel-spectrograms, and further converted to waveform by a pretrained mel vocoder. Compared to training a token-to-wav vocoder in an end-to-end fashion, training a token-to-mel model is more convenient and does not need adversarial training. To better control timbre, a mask strategy is introduced into the training process where the model only computes loss on the un-masked part of spectrograms [du2024cosyvoice]. During inference, spectrogram from speaker prompt conditions the generative process, which can be regarded as a form of “in-context learning”. However, this requires tokens to be extracted from reference prompts before synthesis. Also, inference efficiency may be compromised for better generation quality with multiple inference steps, and this method is only validated on massive amount of data currently.
V Length Reduction and Variable-Rate Tokenization
@@
In most cases, the frame rate of discrete speech tokens ranges from 25 to 100Hz. This leads to a huge discrepancy in lengths between speech representations and the underlying text modality. This discrepancy has been a critical issue in building decoder-only TTS and other LM-based speech generation tasks [wang2024whyspeech], since longer sequences result in harder training and more unstable inference. Therefore, researchers have proposed different ways to mitigate this issue, either by post-processing on tokens, or learning-based variable-rate tokenization.
V-A Length Reduction by Deduplication and Acoustic BPE
@@
Post-training length reduction methods for speech tokens are inspired by language processing techniques. A common approach to reduce token sequence lengths is deduplication [chang23b_interspeech, chang2024exploring], i.e. removing the repeated consecutive tokens in a sequence. Since the encoded continuous features are often close in consecutive frames where the speech dynamics do not change rapidly, they are likely to be quantized to the same unit. Therefore, removing these redundant tokens can yield a more phonetic representation. When the deduplicated tokens are used for generative modeling, a unit-to-speech model (similar to TTS) should be employed to upsample the tokens and convert them back to acoustic signals [lakhotia2021generative].
@@
Another popular approach is acoustic byte-pair encoding (BPE) [^4] The term “acoustic” here is used to distinguish it from traditional BPE applied to text tokens, rather than referring to “acoustic tokens”. or so-called subword modeling [hayashi2020discretalk, ren22_interspeech, chang23b_interspeech, shen2024acoustic, dekel24_interspeech]. Similar to text BPE [Gage1994ANA], acoustic BPE iteratively merges the two most frequent consecutive tokens and adds the merged token to the vocabulary. After training on a corpus, a deterministic BPE mapping is established between original token combinations and the new vocabulary. This mapping enables a lossless compression algorithm, allowing tokens to be perfectly reconstructed after BPE decoding. This operation can identify certain morphological patterns in token sequences, and offers a powerful way to remove redundant tokens. In practice, acoustic BPEs on HuBERT semantic tokens has demonstrated significant speed and performance gains in ASR [chang23b_interspeech, chang2024exploring], spoken language modeling [shen2024acoustic, dekel24_interspeech] and TTS [li24qa_interspeech, vectokspeech].
@@
Although deduplication is a simple and training-free method, acoustic BPE offers several unique advantages over it. First, acoustic BPE can identify redundant patterns that are not simply repetitions, whereas deduplication only removes exact duplicates. Second, deduplication discards the duration information of every token in the resulting sequence. This could be problematic for downstream tasks, as important rhythmic information may reside in the repetitions of tokens. In contrast, acoustic BPE preserves duration information by encoding repetitions of varying lengths into distinct new tokens. Third, acoustic BPE is more flexible than deduplication in terms of target vocabulary size, which can be adjusted based on the desired length reduction ratio and downstream performance.
@@
We visualize the length reduction effect of BPE on different speech tokens in Fig.8, including acoustic and semantic tokens all with a single codebook. From Fig.8, it is evident that different types of tokens exhibit very distinct patterns. Semantic tokens generally show significant length reduction when applying BPE. For single-codebook acoustic tokens, speaker-decoupled LSCodec tokens shows more reduction than general-purpose WavTokenizer and BigCodec. This suggest that the effect of BPE is negatively correlated with the information density in the speech tokens: the less information, the more length reduction achieved by BPE.
V-B Variable Frame Rate Tokens and Unit Discovery